# EXAONE 3.5: Series of Large Language Models for Real-world Use Cases

### v2

@@ 1 Introduction
EXAONE 3.0 instruction-tuned large language model with 7.8B parameters [41] demonstrated strong bilingual capabilities in Korean and English with exceptional real-world performance and instruction-following proficiency. Since its release, we have received diverse feedback from both academic and industrial communities. For instance, academic researchers have emphasized the need for smaller models that can be trained and deployed on low-specification GPUs due to limited access to advanced computational infrastructure. The industry has expressed strong demand for larger models with enhanced performance that remain cost-effective, as well as smaller models suitable for on-device deployment. Additionally, with the increasing adoption of retrieval-augmented generation (RAG) techniques, which generate answers based on reference documents or web search results, there has been substantial demand for models capable of effectively handling longer contexts.
@@ 2.1 Model Configurations
The EXAONE 3.5 language models are based on the latest decoder-only Transformer architecture, and detailed configurations are described in Table 1. These models are identical in structure to the EXAONE 3.0 7.8B model but mainly differ in their configurations related to sizes. Notably, the EXAONE 3.5 language models extend the maximum context length from 4,096 tokens in EXAONE 3.0 to 32,768 tokens by adopting the long-context fine-tuning [7]. All three models share the same vocabulary, which consists roughly of 50% Korean and 50% English.
@@
Non-linearity SwiGLU [44] SwiGLU SwiGLU
@@
Head type GQA [3] GQA GQA
@@
RoPE theta [46] 1,000,000 1,000,000 1,000,000
Tokenizer BBPE [51] BBPE BBPE
@@ 2.2.1 Context Length Extension
To extend the context length, we utilize the long-context fine-tuning technique [7]. To mitigate the catastrophic forgetting problem [32], where the model forgets what it learned during the first pre-training stage, a replay-based method [2] is applied. Specifically, during the second-stage pre-training, we reuse a portion of the data used in the first-stage. While documents exceeding the maximum context length is split into smaller chunks in the first-stage, the original corpus are trained without being divided into chunks in the second-stage to extend the models’ context length.
@@ 2.2.2 Decontamination
By the nature of massively web-crawled corpus, test-set examples often appear in the training corpus [43, 55]. These contaminated examples are likely to harm generalization performance and confuse test metrics, thus presenting unfair evaluations to users. To prevent the contaminated examples undermine the generalization performance of EXAONE 3.5 language models, we rigorously apply a decontamination process for all targeted benchmark test data and remove contaminated examples from the training pipeline.
We borrow a simple yet powerful substring-level matching method [36] with stricter criteria. The entire decontamination process is described in Figure 4 in Appendix C. We first normalize all test-set examples by removing all other characters except alphabets and numbers, then we extract all unique substrings with sliding window size $S=50$ and a stride of 1. To determine whether a training example is contaminated, we randomly sample $N=10$ substrings from the normalized training example and check if they exist in the substring pools. Table C in Appendix C provides examples of documents found in web corpora considered as contaminated.
@@ 2.2.3 Training Cost
Considering the computational cost of pre-training a large language model (LLM), it is necessary to make the training efficient by achieving as much high performance as possible with limited resources. Table 3 compares the total amounts of computations required for pre-training between the EXAONE 3.5 32B language model and others of similar size. When we simply approximate the total amounts of computations as the product of the model size and the number of training tokens [24, 19], Qwen 2.5 32B, for example, requires 2.77 times more computations than EXAONE 3.5 32B. One of the noticeable characteristics of the EXAONE 3.5 language models is that they demonstrate high performance despite being trained at lower costs than the other baseline models (see Section 3).
@@ 2.3 Post-training
After pre-training, models go through further processes for strengthening their instruction-following capabilities and aligning with human preferences, which are well known as supervised fine-tuning (SFT) [53] and preference optimization.
@@ 2.3.1 Supervised Fine-tuning
To perform well on new or unseen instructions, a model needs to be trained on pairs of instruction-response datasets with varying difficulty from different domains. Hence, in order to build training data covering a wide range of fields, we extract core knowledge from 8M web corpora using a taxonomic system, as shown in Figure 1. We then generate an instruction-tuning dataset based on the extracted knowledge taxonomy. Finally, leveraging an instruction evolution method, which stems from the method proposed in [58], we diversify the complexity levels so that instructions with various complexities and difficulties can be produced.
Figure 1: A procedure of instruction-tuning data construction. First, we extract the core knowledge from large-volume web corpora and classify it within the taxonomy we defined in advance. Next, instruction-tuning data is generated based on the knowledge. To construct additional training data that is more complex, we leverage an instruction-evolving method [58] that lets the final dataset cover various fields with varying levels of difficulty.
@@ 2.3.2 Preference Optimization
Direct alignment algorithms (DAAs) [38], such as DPO [39] and SimPO [33], are used to train models after supervised fine-tuning to align models with human preferences. We create preference data for training using synthetic data and pre-collected data. For response generation, we sample $N$ responses from multiple models for the prompt $x$ drawn from the preference data and select the best response as $y_{w}$ and the worst response as $y_{l}$ based on the scores of a reward model to create a preference data, $\{x,y_{w},y_{l}\}$. To validate preference data, we use an additional reward model to calculate agreement based on the rankings of the two reward models and filter out data with agreement below the threshold. Our preference optimization comprises multiple stages to sequentially train models $M_{1}$ and $M_{2}$ through DAAs, where $M_{0}$ is initialized from the SFT model. The staged pipeline enables us to mitigate over-optimization [38] that may occur during the DAAs’ training process. Figure 2 shows a schematic diagram for constructing our preference dataset and training process.
@@ 2.4 Data Compliance
Developing AI models requires a large amount of data, and the acquisition and utilization of this data can lead to various legal issues, such as copyright infringement, intellectual property infringement, and personal information protection violations. To minimize these risks, LG AI Research conducts AI Compliance reviews throughout the entire process of data collection, AI model training, and information provision. For more detailed information, please refer to the EXAONE 3.0 Technical Report [41] and the LG AI Ethics Principles [28].
@@ 3 Evaluation
This section presents the evaluation settings and results of EXAONE 3.5 language models on various benchmark datasets. We select recently released open language models for baselines of our models to compare our performances on the benchmarks. All baselines and their detailed information are described in Appendix D.1.
@@ 3.1 Benchmarks
Real-world Use Cases MT-Bench [59] EN LLM-as-a-judge (judge: gpt-4o-2024-08-06) [^0] The separability of the original GPT-4 judge results is notably low, prompting the adoption of gpt-4o-2024-08-06 as judge. LLM score
LiveBench [54] (v2024-08-31) EN Ground-truth match Accuracy
Arena-Hard-v0.1 [29] EN LLM-as-a-judge (judge: gpt-4-1106-preview) Win rate
AlpacaEval 2.0 LC [12] EN LLM-as-a-judge (judge: gpt-4-1106-preview) Win rate
IFEval[61] EN Prompt-level / strict accuracy Accuracy
KoMT-Bench [42] KO LLM-as-a-judge (judge: gpt-4o-2024-08-06) LLM score
LogicKor [37] KO LLM-as-a-judge (judge: gpt-4-1106-preview) LLM score
Long Context Needle-In-A-Haystack [23] EN/KO Ground-truth match Accuracy
LongBench [5] EN Ground-truth match F1, Rouge
LongRAG [21] (extended) EN LLM-as-a-judge (judge: gpt-4o-2024-08-06) LLM score
@@
General Domain GSM8K [9] EN 0-shot / CoT Accuracy
MATH [17, 27] EN 0-shot / CoT Accuracy
HumanEval [6] EN 0-shot pass@1
MBPP [4] EN 0-shot (Evalplus base) [^1] We choose the MBPP base from EvalPlus [31], which is a subset of the original and consists of refined, high-quality problems. pass@1
GPQA [40] EN 0-shot / CoT Accuracy
ARC-C [8] EN 0-shot Accuracy
BBH [47] EN 0-shot / CoT Accuracy
MMLU [16] EN 0-shot / CoT Accuracy
KMMLU [45] KO 0-shot / CoT Accuracy
@@ 3.2 Overall Performance
The results of overall performance against three categories are presented in Table 5. Our EXAONE 3.5 language models, with sizes 32B and 7.8B, perform best in Real-world Use Cases and Long Context categories compared to baseline models while showing competitive results in the General Domain category. Our smallest model, EXAONE 3.5 2.4B, outperforms baselines with similar sizes in all three categories, demonstrating strong performance. Surprisingly, our 2.4B model, despite its small size, has shown better performance compared to baselines even with a larger size (< 9B) except for Qwen 2.5 7B in General Domain. Considering the recent surge in demand for smaller large language models (sLLM) [52], we believe that our EXAONE 3.5 2.4B model is well-positioned to be highly competitive in both academic and industrial use.
@@
Qwen 2.5 32B [49] 69.8 66.9 78.7
C4AI Command R 32B  [10] 46.0 63.4 56.8
Gemma 2 27B [48] 64.2 - 68.7
Yi 1.5 34B [2] 46.9 - 53.9
@@
Qwen 2.5 7B [49] 52.7 56.1 71.0
Llama 3.1 8B [15] 48.6 58.8 62.4
Gemma 2 9B [48] 57.9 - 62.9
Phi 3 small (7B) [1] 41.7 33.4 63.2
@@
Qwen 2.5 3B [49] 44.5 40.7 62.1
Qwen 2.5 1.5B [49] 30.1 34.5 47.9
Llama 3.2 3B [34] 36.7 44.2 54.9
Gemma 2 2B [48] 41.7 - 42.2
Table 5: Overall comparison results of EXAONE 3.5 language models with similar-sized baseline language models. Here, a dash (-) indicates the model does not support context lengths longer than 16K. Bold scores indicate the best performance, and underlined scores mean the second best. The detailed information for each baseline is described in Appendix D.1.
@@ 3.4.1 Needle-in-a-Haystack
Needle-in-a-Haystack (NIAH) [23] serves as a benchmark to assess how effectively models can locate and retrieve information hidden at random locations within long documents. We comprehensively evaluate our models’ ability to process and retrieve information from long contexts, up to 32K tokens. Furthermore, we extend NIAH to Korean and employ it to evaluate our models’ long context processing ability across both English and Korean contexts.
@@ 3.4.2 Long Context Understanding
To assess long context understanding capabilities, we evaluate our models using benchmarks including LongBench [5] and LongRAG [21]. We expand unanswerable cases in LongRAG to make it more challenging. We also build Ko-LongRAG, the Korean counterpart to LongRAG, to evaluate long context understanding in Korean. For more realistic RAG scenario, requiring answers to difficult questions using actual web-searched results, we constructed Ko-WebRAG benchmark. We refer readers to the Appendix D.2 for more details.
@@
Table 7: Performance comparison results of EXAONE 3.5 language models with similar-sized recently released language models across four benchmarks representing long context scenarios. A dash (-) indicates that the model does not support context lengths longer than 16K. Context lengths for each model are detailed in Table 11. The average score in the rightmost is calculated as a macro average across the benchmarks. Bold scores indicate the best performance, and underlined scores mean the second best.
@@ 3.5 General Domain
To better simulate the real-world scenarios where a chatbot model usually receives a single query from users, we evaluate all benchmarks in the General Domain category using the 0-shot setting. To achieve this, we prompt language models with instructions that require specific answer formats and parse the final answer from the responses. For a fair comparison, we use the same prompts across all models. We make public all the prompts we used in Appendix D.3 for transparent reproducibility.
@@ 4 Responsible AI
EXAONE 3.5 language models were developed in accordance with the Responsible AI Development Framework encompassing data governance, ethical considerations, and risk management as it would be made available to a wide range of users. Given the nature of open models – eventually leading to wide use in various domains – we aim to maximize social benefits while ensuring humanity, fairness, safety, accountability, and transparency as mandated by the LG AI Ethics Principles [28].

### v3

@@ 1 Introduction
EXAONE 3.0 instruction-tuned large language model with 7.8B parameters [research2024exaone3078binstruction] demonstrated strong bilingual capabilities in Korean and English with exceptional real-world performance and instruction-following proficiency. Since its release, we have received diverse feedback from both academic and industrial communities. For instance, academic researchers have emphasized the need for smaller models that can be trained and deployed on low-specification GPUs due to limited access to advanced computational infrastructure. The industry has expressed strong demand for larger models with enhanced performance that remain cost-effective, as well as smaller models suitable for on-device deployment. Additionally, with the increasing adoption of retrieval-augmented generation (RAG) techniques, which generate answers based on reference documents or web search results, there has been substantial demand for models capable of effectively handling longer contexts.
@@ 2.1 Model Configurations
The EXAONE 3.5 language models are based on the latest decoder-only Transformer architecture, and detailed configurations are described in Table 1. These models are identical in structure to the EXAONE 3.0 7.8B model but mainly differ in their configurations related to sizes. Notably, the EXAONE 3.5 language models extend the maximum context length from 4,096 tokens in EXAONE 3.0 to 32,768 tokens by adopting the long-context fine-tuning [chen2023extendingcontextwindowlarge]. All three models share the same vocabulary, which consists roughly of 50% Korean and 50% English.
@@
Non-linearity SwiGLU [shazeer2020gluvariantsimprovetransformer] SwiGLU SwiGLU
@@
Head type GQA [ainslie2023gqatraininggeneralizedmultiquery] GQA GQA
@@
RoPE theta [su2023roformerenhancedtransformerrotary] 1,000,000 1,000,000 1,000,000
Tokenizer BBPE [wang2020bbpe] BBPE BBPE
@@ 2.2.1 Context Length Extension
To extend the context length, we utilize the long-context fine-tuning technique [chen2023extendingcontextwindowlarge]. To mitigate the catastrophic forgetting problem [MCCLOSKEY1989109], where the model forgets what it learned during the first pre-training stage, a replay-based method [ai2024yiopenfoundationmodels] is applied. Specifically, during the second-stage pre-training, we reuse a portion of the data used in the first-stage. While documents exceeding the maximum context length is split into smaller chunks in the first-stage, the original corpus are trained without being divided into chunks in the second-stage to extend the models’ context length.
@@ 2.2.2 Decontamination
By the nature of massively web-crawled corpus, test-set examples often appear in the training corpus [riddell2024quantifying, xu2024benchmarking]. These contaminated examples are likely to harm generalization performance and confuse test metrics, thus presenting unfair evaluations to users. To prevent the contaminated examples undermine the generalization performance of EXAONE 3.5 language models, we rigorously apply a decontamination process for all targeted benchmark test data and remove contaminated examples from the training pipeline.
We borrow a simple yet powerful substring-level matching method [openai2023GPT4technicalreport] with stricter criteria. The entire decontamination process is described in Figure 4 in Appendix C. We first normalize all test-set examples by removing all other characters except alphabets and numbers, then we extract all unique substrings with sliding window size $S=50$ and a stride of 1. To determine whether a training example is contaminated, we randomly sample $N=10$ substrings from the normalized training example and check if they exist in the substring pools. Table LABEL:tab:decon_example in Appendix C provides examples of documents found in web corpora considered as contaminated.
@@ 2.2.3 Training Cost
Considering the computational cost of pre-training a large language model (LLM), it is necessary to make the training efficient by achieving as much high performance as possible with limited resources. Table 3 compares the total amounts of computations required for pre-training between the EXAONE 3.5 32B language model and others of similar size. When we simply approximate the total amounts of computations as the product of the model size and the number of training tokens [kaplan2020scalinglawsneurallanguage, hoffmann2022an], Qwen 2.5 32B, for example, requires 2.77 times more computations than EXAONE 3.5 32B. One of the noticeable characteristics of the EXAONE 3.5 language models is that they demonstrate high performance despite being trained at lower costs than the other baseline models (see Section 3).
@@ 2.3 Post-training
After pre-training, models go through further processes for strengthening their instruction-following capabilities and aligning with human preferences, which are well known as supervised fine-tuning (SFT) [wei2022finetunedlanguagemodelszeroshot] and preference optimization.
@@ 2.3.1 Supervised Fine-tuning
To perform well on new or unseen instructions, a model needs to be trained on pairs of instruction-response datasets with varying difficulty from different domains. Hence, in order to build training data covering a wide range of fields, we extract core knowledge from 8M web corpora using a taxonomic system, as shown in Figure 1. We then generate an instruction-tuning dataset based on the extracted knowledge taxonomy. Finally, leveraging an instruction evolution method, which stems from the method proposed in [zeng-etal-2024-automatic], we diversify the complexity levels so that instructions with various complexities and difficulties can be produced.
Figure 1: A procedure of instruction-tuning data construction. First, we extract the core knowledge from large-volume web corpora and classify it within the taxonomy we defined in advance. Next, instruction-tuning data is generated based on the knowledge. To construct additional training data that is more complex, we leverage an instruction-evolving method [zeng-etal-2024-automatic] that lets the final dataset cover various fields with varying levels of difficulty.
@@ 2.3.2 Preference Optimization
Direct alignment algorithms (DAAs) [rafailov2024scaling], such as DPO [rafailov2023directpreferenceoptimizationlanguage] and SimPO [meng2024simposimplepreferenceoptimization], are used to train models after supervised fine-tuning to align models with human preferences. We create preference data for training using synthetic data and pre-collected data. For response generation, we sample $N$ responses from multiple models for the prompt $x$ drawn from the preference data and select the best response as $y_{w}$ and the worst response as $y_{l}$ based on the scores of a reward model to create a preference data, $\{x,y_{w},y_{l}\}$. To validate preference data, we use an additional reward model to calculate agreement based on the rankings of the two reward models and filter out data with agreement below the threshold. Our preference optimization comprises multiple stages to sequentially train models $M_{1}$ and $M_{2}$ through DAAs, where $M_{0}$ is initialized from the SFT model. The staged pipeline enables us to mitigate over-optimization [rafailov2024scaling] that may occur during the DAAs’ training process. Figure 2 shows a schematic diagram for constructing our preference dataset and training process.
@@ 2.4 Data Compliance
Developing AI models requires a large amount of data, and the acquisition and utilization of this data can lead to various legal issues, such as copyright infringement, intellectual property infringement, and personal information protection violations. To minimize these risks, LG AI Research conducts AI Compliance reviews throughout the entire process of data collection, AI model training, and information provision. For more detailed information, please refer to the EXAONE 3.0 Technical Report [research2024exaone3078binstruction] and the LG AI Ethics Principles [lgethics].
@@ 3 Evaluation
This section presents the evaluation settings and results of EXAONE 3.5 language models on various benchmark datasets. We select recently released open language models for baselines of our models to compare our performances on the benchmarks. All baselines and their detailed information are described in Appendix LABEL:appendix:baseline_details.
@@ 3.1 Benchmarks
Real-world Use Cases MT-Bench [zheng2023judging] EN LLM-as-a-judge (judge: gpt-4o-2024-08-06) [^0] The separability of the original GPT-4 judge results is notably low, prompting the adoption of gpt-4o-2024-08-06 as judge. LLM score
LiveBench [white2024livebenchchallengingcontaminationfreellm] (v2024-08-31) EN Ground-truth match Accuracy
Arena-Hard-v0.1 [li2024crowdsourceddatahighqualitybenchmarks] EN LLM-as-a-judge (judge: gpt-4-1106-preview) Win rate
AlpacaEval 2.0 LC [dubois2024lengthcontrolledalpacaevalsimpleway] EN LLM-as-a-judge (judge: gpt-4-1106-preview) Win rate
IFEval[zhou2023instructionfollowingevaluationlargelanguage] EN Prompt-level / strict accuracy Accuracy
KoMT-Bench [komt-bench] KO LLM-as-a-judge (judge: gpt-4o-2024-08-06) LLM score
LogicKor [logickor] KO LLM-as-a-judge (judge: gpt-4-1106-preview) LLM score
Long Context Needle-In-A-Haystack [kamradt2023llmtest] EN/KO Ground-truth match Accuracy
LongBench [bai-etal-2024-longbench] EN Ground-truth match F1, Rouge
LongRAG [jiang2024longragenhancingretrievalaugmentedgeneration] (extended) EN LLM-as-a-judge (judge: gpt-4o-2024-08-06) LLM score
@@
General Domain GSM8K [cobbe2021trainingverifierssolvemath] EN 0-shot / CoT Accuracy
MATH [hendrycks2021measuringmathematicalproblemsolving, minerva_math] EN 0-shot / CoT Accuracy
HumanEval [chen2021evaluatinglargelanguagemodels] EN 0-shot pass@1
MBPP [austin2021programsynthesislargelanguage] EN 0-shot (Evalplus base) [^1] We choose the MBPP base from EvalPlus [liu2023is], which is a subset of the original and consists of refined, high-quality problems. pass@1
GPQA [rein2023gpqagraduatelevelgoogleproofqa] EN 0-shot / CoT Accuracy
ARC-C [clark2018thinksolvedquestionanswering] EN 0-shot Accuracy
BBH [suzgun2022challenging] EN 0-shot / CoT Accuracy
MMLU [hendrycks2020measuring] EN 0-shot / CoT Accuracy
KMMLU [son2024kmmlumeasuringmassivemultitask] KO 0-shot / CoT Accuracy
@@ 3.2 Overall Performance
The results of overall performance against three categories are presented in Table 5. Our EXAONE 3.5 language models, with sizes 32B and 7.8B, perform best in Real-world Use Cases and Long Context categories compared to baseline models while showing competitive results in the General Domain category. Our smallest model, EXAONE 3.5 2.4B, outperforms baselines with similar sizes in all three categories, demonstrating strong performance. Surprisingly, our 2.4B model, despite its small size, has shown better performance compared to baselines even with a larger size (< 9B) except for Qwen 2.5 7B in General Domain. Considering the recent surge in demand for smaller large language models (sLLM) [sllmsurvey2024], we believe that our EXAONE 3.5 2.4B model is well-positioned to be highly competitive in both academic and industrial use.
@@
Qwen 2.5 32B [qwen2.5] 69.8 66.9 78.7
C4AI Command R 32B  [cohere_for_ai_2024] 46.0 63.4 56.8
Gemma 2 27B [gemmateam2024gemma2improvingopen] 64.2 - 68.7
Yi 1.5 34B [ai2024yiopenfoundationmodels] 46.9 - 53.9
@@
Qwen 2.5 7B [qwen2.5] 52.7 56.1 71.0
Llama 3.1 8B [grattafiori2024llama3herdmodels] 48.6 58.8 62.4
Gemma 2 9B [gemmateam2024gemma2improvingopen] 57.9 - 62.9
Phi 3 small (7B) [abdin2024phi3technicalreporthighly] 41.7 33.4 63.2
@@
Qwen 2.5 3B [qwen2.5] 44.5 40.7 62.1
Qwen 2.5 1.5B [qwen2.5] 30.1 34.5 47.9
Llama 3.2 3B [llama3.2] 36.7 44.2 54.9
Gemma 2 2B [gemmateam2024gemma2improvingopen] 41.7 - 42.2
Table 5: Overall comparison results of EXAONE 3.5 language models with similar-sized baseline language models. Here, a dash (-) indicates the model does not support context lengths longer than 16K. Bold scores indicate the best performance, and underlined scores mean the second best. The detailed information for each baseline is described in Appendix LABEL:appendix:baseline_details.
@@ 3.4.1 Needle-in-a-Haystack
Needle-in-a-Haystack (NIAH) [kamradt2023llmtest] serves as a benchmark to assess how effectively models can locate and retrieve information hidden at random locations within long documents. We comprehensively evaluate our models’ ability to process and retrieve information from long contexts, up to 32K tokens. Furthermore, we extend NIAH to Korean and employ it to evaluate our models’ long context processing ability across both English and Korean contexts.
@@ 3.4.2 Long Context Understanding
To assess long context understanding capabilities, we evaluate our models using benchmarks including LongBench [bai-etal-2024-longbench] and LongRAG [jiang2024longragenhancingretrievalaugmentedgeneration]. We expand unanswerable cases in LongRAG to make it more challenging. We also build Ko-LongRAG, the Korean counterpart to LongRAG, to evaluate long context understanding in Korean. For more realistic RAG scenario, requiring answers to difficult questions using actual web-searched results, we constructed Ko-WebRAG benchmark. We refer readers to the Appendix LABEL:appendix:lc_setting for more details.
@@
Table 7: Performance comparison results of EXAONE 3.5 language models with similar-sized recently released language models across four benchmarks representing long context scenarios. A dash (-) indicates that the model does not support context lengths longer than 16K. Context lengths for each model are detailed in Table LABEL:tab:baselines. The average score in the rightmost is calculated as a macro average across the benchmarks. Bold scores indicate the best performance, and underlined scores mean the second best.
@@ 3.5 General Domain
To better simulate the real-world scenarios where a chatbot model usually receives a single query from users, we evaluate all benchmarks in the General Domain category using the 0-shot setting. To achieve this, we prompt language models with instructions that require specific answer formats and parse the final answer from the responses. For a fair comparison, we use the same prompts across all models. We make public all the prompts we used in Appendix LABEL:appendix:general_dom_setting for transparent reproducibility.
@@ 4 Responsible AI
EXAONE 3.5 language models were developed in accordance with the Responsible AI Development Framework encompassing data governance, ethical considerations, and risk management as it would be made available to a wide range of users. Given the nature of open models – eventually leading to wide use in various domains – we aim to maximize social benefits while ensuring humanity, fairness, safety, accountability, and transparency as mandated by the LG AI Ethics Principles [lgethics].